OpenStudy (anonymous):

How would I change this rsin(theat)=3 into rectangular coordinate, do I just solve for r and then plug it into to y=rsin(theta) and x=rcos(theta)?

OpenStudy (anonymous):

Start by multiplying both sides of rsin(theta)=3 by r

OpenStudy (anonymous):

\[rsin(\theta)=3 \] \[r^2\sin(\theta)=3r\] but \[r^2=x^2+y^2\] and \[r=\sqrt(x^2+y^2)\] and \[\sin(\theta)=y/r\] So \[r^2\sin(\theta)=3r \] BECOMES \[(x^2+y^2)(y/\sqrt(x^2+y^2))=3\sqrt(x^2+y^2)\] Simplify and this becomes, y=3

OpenStudy (beginnersmind):

krystabell's answer is nice, but I think the simplest solution is in your question. "How would I change this ****rsin(theat)=3**** into rectangular coordinate, do I just solve for r and then plug it into to ****y=rsin(theta)**** and x=rcos(theta)? "

OpenStudy (beginnersmind):

For the record I only saw it after I read through krystabell's solution :)

OpenStudy (anonymous):

Hehehe yes oh those formula's they sure do make you do what you did in the last problems...i didn't address that. so you don't want to solve directly for r because you would be left with r=3/sin(theta) and this would get you nowhere...Thanks beginnersmind for pointing this out!!!